mathslover (mathslover):

What is the magnitude of gravitational force exerted by a 15 kg mass on a25 kg mass separated by a distance of 25 Cm? What is the acceleration produced on each mass?

mathslover (mathslover):

I did like this : \[\large{\color{red}{F=\frac{Gm_1m_2}{R^2}}}\] \[\large{\color{blue}{F=\frac{6.7*10^{-11}*15 kg*25 kg}{0.25*0.25}}}\] \[\large{\color{green}{F=\frac{6.7*15*25*10^{-7}}{25*25}}}\] \[\large{\color{brown}{F=\frac{6.7*15*10^{-7}}{25}}}\] \[\large{\color{orange}{F=\frac{6.7*3*10^{-7}}{5}}}\] \[\large{\color{magenta}{F={4*10^{-7}N\implies \textbf{Approx.}}}}\]

mathslover (mathslover):

@lgbasallote @rajathsbhat @experimentX @Callisto @hartnn

OpenStudy (anonymous):

u r absolutely right! :)

mathslover (mathslover):

now : F = m * a hence for m = 15 kg : \[\large{\color{red}{4*10^{-7}=15 kg * a}}\] \[\large{\color{violet}{\frac{4*10^{-7}}{15}ms^{-2}=a}}\] Now for m = 25 kg \[\large{\color{blue}{4*10^{-7}=25 kg*a}}\] \[\large{\color{green}{\frac{4*10^{-7}}{25}ms^{-2}=a}}\]

OpenStudy (anonymous):

Right again! So....... what is ur doubt?

mathslover (mathslover):

m i right? that is my doubt

OpenStudy (anonymous):

my medal says it all!

OpenStudy (lgbasallote):

this is making me epileptic :/

mathslover (mathslover):

thanks rajath and lgba for medal

OpenStudy (anonymous):

yw :D

mathslover (mathslover):

@sauravshakya @sami-21 can u confirm me this again .. ?

mathslover (mathslover):

actually this is not much complex as per the level of my book .. I thought the answer must be integer type..

mathslover (mathslover):

@heena

mathslover (mathslover):

@ajprincess

mathslover (mathslover):

m i right heena?

OpenStudy (anonymous):

yep and here u knw force is taken same for both mass becoz both aplly same amount of force on each other ok

mathslover (mathslover):

right thanks a lot heena @waterineyes

OpenStudy (ujjwal):

That will be initial acceleration. And yeah, the acceleration on each body will increase gradually.. As, the bodies move towards each other, the gravitational force between them will increase and thus acceleration will also increase gradually. In actual practice, that effect may not be felt much due to less masses but in theory, it will still exist!

mathslover (mathslover):

thanks a lot @ujjwal

OpenStudy (anonymous):

wow..u got a lot of confirmations for ur answer! here's mine: yup, u're absolutely right ;)

mathslover (mathslover):

thanks a lot.. actually this is very imp. question for me for exams hence need that confirmation

OpenStudy (shane_b):

You got 6 medals for answering your own question. Here, I will give you another since you got it right :)

mathslover (mathslover):

"you got it right :)" thanks for this

OpenStudy (kainui):

You got it completely wrong.

OpenStudy (kainui):

Just kidding. =P

OpenStudy (shane_b):

And I thought I missed something...I looked at it pretty fast and moved on :)

OpenStudy (anonymous):

What a colourful working .. and you got it right too ;)

mathslover (mathslover):

thanks @kainui and @supercrazy92